the vertical cross section of a 40 foot high masonry dam is uniform. The dam is 80 feet long and weighs 125 pounds per cubic foot. Determine the dam's weight


1
Expert's answer
2021-08-17T16:52:39-0400

Dam's weight=weight of material*volume of dam

Volume of dam=c.a*length

The question as it is given isn't complete. The complete question provides that the width and breadth of the dam is 4ft and 16ft respectively.

With this ;

="\frac{1}{2}(a+b)h*l"

="\frac{1}{2}(4+16)40*80"

=32000 lb3

Dam's weight=32000*125

=4,000,000 lbs.